MULTICHANNEL QUANTUM DEFECT THEORY OF THE FORMYL RADICAL

\underline{E.J. ZUCKERMAN}, Department of Chemistry and Physics, Augusta State University, Augusta, GA 30904; E.R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907; R.P. BRINT, Department of Chemistry, National University of Ireland, Cork, Ireland; C. JUNGEN, Laboratoire de Photophysique Molecular du CNRS, Universite Paris Sud, Orsay, France.

The high Rydberg states of the formyl radical (HCO) have been examined in great detail by means of double-resonance spectroscopy of vibrationally autoionizing states. Simulations utilizing the Rydberg formula have described these states in terms of a phenomological quantum defect (δ). We have used these assignments as a guide in applying the multichannel quantum defect theory (MQDT) to HCO. The current calculations include intensity as well as energy considerations, and will serve as a further guide toward the development of a complete polyatomic MQDT formalism.